17. At 30 °C Kw = 1.471 x 10 -14. What is the [ H] and [OH-] at a neutral pH? ( 12 pt)
Answer #1

Answer. OH = 44x) + OH Pay at 30° Kw = [ht] [H-] ato neutral pH +] = (04-) = 2 Kw = (1) (n) 1.471810-14 = 22 r = 1.21 X.10-7